System Software Teleconference Meeting Minutes - 2007-08-21

Mitch Bradley wmb at
Tue Aug 21 22:19:59 EDT 2007

OLPC System Software Teleconference, 2007-08-21

Participants: Jim, Mitch, Jordan, Chris Ball, Quanta, Andres, wad, J5, 
Noah, kim, Joel, richard

== Status ==

Action items from last week:

#1396 -- We're waiting on the problem with keyboard repeat events.
* AI: Richard to get us a release firmware with this patched.
The "goto hack" made it in the v24 EC firmware.  We think this
fixes the repeat problem.

#1500 -- Need public key for mfg-data.
* AI: Jim to talk to Noah/Ivan to understand why we still don't
      have the public keys generated and signing working yet.

We have a key format - Noah has code that "documents" the format.
The problem is getting the hash functions installed in the HSM machine.
Noah to supply to wmb the key format, a key, a signature file, and
a signed object.

#1547 -- Decide frame buffer depth.
* AI: cjb/bernie to performance-test color depth change.
No progress yet.

Bernie: evdev stopped working.  Before switching to 1.7, we were using
the respeclaration instead of the HAL.
We could switch to Xorg 1.4 and shoehorn in the old keyboard/mouse driver.
Jg just wants some packages that we can use for testing, but they
don't have to go into builds.

80 fps in 24 bit color, 50 fps in 16-bit color

cjb: the 16-bit code is not using MMX, possible further so we might want

#1748 -- USB key upgrades currently disabled in USB autoreinstall
* AI: cscott to enable them.
+ Done, no complaints so far.

#2184 -- OHM idleness detection isn't working
* AI: jg and cjb to meet and decide what OHM should do for Trial-3,
      mail the result to Richard Hughes and devel at .
+ Spec mailed out, but no final resolution yet.  We need to get
the results into the wiki (AI jg).

#2237 -- Need to find the best DCON write combination for power-savings
* AI: Joel to do so and pass on to Andres to implement.
Joel did the first part but not the second part


cjb has a proposed fixed firmware for the Marvel hang, has already
passed 200 resumes, albeit in mesh mode.  If and when it does fail,
it will spit out a bunch of interesting debug information.

Builds are messed up right now because of activity problems (color module
went away).  People need to remove dependencies on modules that they 
don't use.
TamTam is having an issue with alpha and other startup problems.
Paint needs a few changes.  Hippi (python console app by cjb) is working
well (the icon needs vetting).  The icon widths need fixing.

Running measure still breaks the sound.

  200 USB-to-XOserial adapters are on order.
  has a couple of extra old dongles
  has 20 cables
  Quanta has 80 cables in inventory and is building another 120.

School server status

New pango released with support for Ethopian languages.

Blocker bug triage:

#1339 - proposed patch available, but not verified

#1396 - believed fixed, needs to be retested

#1431 - pending schoolserver work - reassign to wad

#1496 -
#1500 - pending keys

#1835 - reassign to cjb

We don't have the Ethiopian keyboard working - missing glyphs, compose 
key problem, and others.

#2182 - mostly fixed in Ctest, but you can always find a battery that is 
at just the right
voltage to trigger it.  Close the bug.

#2237 - On the list...

#2243 - Needs to be tested with Ctest.

#2449 - close

#2586 - check when we get Ctest systems

#2833 - jg to ping erikb

More information about the Devel mailing list